OpenStudy (anonymous):

Find the equation of tangent to the curve for y=-3x^2+5x+3

OpenStudy (jusaquikie):

what class is this for, calculus?

OpenStudy (anonymous):

yes it is calculus

OpenStudy (jusaquikie):

it's asking for the derivative do you know how to take the derivative yet?

OpenStudy (anonymous):

so the derivative is-6x+5

OpenStudy (jusaquikie):

yes

OpenStudy (anonymous):

i dont know what to do after that

OpenStudy (jusaquikie):

what form do they want the answer in? did they give you a point on the curve?

OpenStudy (anonymous):

the problem says for the curve equation y=-3x^2+5x+3, find equation of tangent to the curve

OpenStudy (anonymous):

part b of the problem is find the slope of the tangent at x=3

OpenStudy (anonymous):

part c of the aproblem is to find the angle of inclination of the tangent

OpenStudy (jusaquikie):

ok the answer to part a is the derivative and for part b put in 3 for x and solve for y

OpenStudy (anonymous):

am i putting the 3 into the original equation or the derived equation

OpenStudy (jusaquikie):

once you have the answer for part b take the arctan of that and it will give you the angle

OpenStudy (jusaquikie):

into the tangent function which is the derivitive

OpenStudy (anonymous):

so its -6(3)+5 correct and three fore i should get a slop of 13.... thta seems like an awfully high slope

OpenStudy (jusaquikie):

y' is the slope at any point x along the curve. so if you want the slope at x=3 replace x with 3 and solve. if tou want the tangent at x=189 put in 189 for x and solve.

OpenStudy (anonymous):

i mean -13

OpenStudy (jusaquikie):

it dosen't matter really, that will equate to -85.6°

OpenStudy (anonymous):

ok thank you

myininaya (myininaya):

If part a is to find the equation of the tangent line, then we aren't totally complete on the answer. You found the formula for the slope of the tangent line at any point on the curve. Linear functions can be written in this form: y-y1=m(x-x1) Say we want to find the tangent line at (a,f(a)) then y=m(x-a)+f(a) Don't forget to replace your m here with y' evaluated at x=a.

OpenStudy (anonymous):

I am confused again...would you be able to explain and show the steps to each part